# Implicit Bias of Per-sample Adam on Separable Data: Departure from the Full-batch Regime

### v2

@@ Appendix E Missing Proofs in Section 3
In this section, we provide the omitted proofs in Section 3. We first introduce the proof of Corollary 3.2 describing how GR datasets eliminate coordinate-adaptivity of Inc-Adam. Then, we review previous literature on the limit direction of weighted GD and prove Theorem 3.3.
@@
Given GR data $\{{\mathbf{x}}_{i}\}_{i\in[N]}$, let $x_{i}=|{\mathbf{x}}_{i}[0]|$. Notice that
@@
We now turn to the proof of Theorem 3.3, building upon the foundational work of Ji et al. [2020], who characterized the convergence direction of GD via its regularization path. Subsequent research has extended this characterization to weighted GD, which optimizes the weighted empirical risk $\mathcal{L}_{\mathbf{q}(t)}({\mathbf{w}})=\sum_{i\in[N]}q_{i}(t)\ell({\mathbf{w}}^{\top}{\mathbf{x}}_{i})$. Xu et al. [2021] proved that weighted GD converges to $\ell_{2}$-max-margin direction on the same linear classification task when the weights are fixed during training. This condition was later relaxed by Zhai et al. [2023], who demonstrated that the same convergence guarantee holds provided the weights converge to a limit, i.e., $\exists\lim_{t\rightarrow\infty}\mathbf{q}(t)=\hat{\mathbf{q}}$.
Our setting, however, introduces distinct technical challenges. First, the weights are bounded but not guaranteed to converge. The most relevant existing result is Theorem 7 in Zhai et al. [2023], which establishes the same limit direction but requires the stronger combined assumptions of lower-bounded weights, loss convergence, and directional convergence of the iterates. A further complication in our analysis is an additional error term, $\bm{\epsilon}_{r}$ in Corollary 3.2, which must be carefully controlled. Our fine-grained analysis overcomes these issues by extending the methodology of Ji et al. [2020], enabling us to manage the error term under the sole, weaker assumption of loss convergence.
@@
By introducing $\bm{a}$}-weighted loss, we can regard weighted GD as vanilla GD with respect to weighted loss. To follow the line of Ji et al. [2020], we show that the regularization path converges in direction to $\ell_{2}$-max-margin solution, regardless of the choice of the weight vector $\bm{a}$} if it is bounded by two positive constants, and such convergence is uniform; we can take sufficiently large $B$ to be close the $\ell_{2}$ solution for any ${\bm{a}}\in[c_{1},c_{2}]^{N}$.
Lemma E.2 (Adaptation of Proposition 10 in Ji et al. [2020]).
@@
We first have to show the uniqueness of $\ell_{2}$-max-margin solution. This proof was introduced by Ji et al. [2020, Proposition 10], but we provide it for completeness. Suppose that there exist two distinct unit vectors ${\mathbf{u}}_{1}$ and ${\mathbf{u}}_{2}$ such that both of them achieve the max-margin $\hat{\gamma}$. Take ${\mathbf{u}}_{3}=\frac{{\mathbf{u}}_{1}+{\mathbf{u}}_{2}}{2}$ as a middle point of ${\mathbf{u}}_{1}$ and ${\mathbf{u}}_{2}$. Then we get
@@
Lemma E.3 (Adaptation of Lemma 9 in Ji et al. [2020]).
@@ Appendix F Missing Proofs in Section 4
which implies $\sum_{t=t_{0}}^{\infty}\eta_{t}\|\nabla\mathcal{L}({\mathbf{w}}_{t})\|_{{\mathbf{P}}^{-1}({\mathbf{w}}_{t})}<\infty$. Since $\sum_{t=t_{0}}^{T}\eta_{t}=\infty$, we get $\|\nabla\mathcal{L}({\mathbf{w}}_{t})\|_{{\mathbf{P}}^{-1}({\mathbf{w}}_{t})}\rightarrow 0$. From (b), we get $\nabla\mathcal{L}({\mathbf{w}}_{t})\rightarrow 0$, and consequently, $\mathcal{L}({\mathbf{w}}_{t})\rightarrow 0$. ∎
@@
Before we provide a rigorous proof of Lemma 4.5, we first demonstrate its intuitive explanation motivated by Soudry et al. [2018]. For simplicity, assume $\ell=\ell_{\exp}$ and let ${\mathbf{w}}_{t}=g(t)\hat{{\mathbf{w}}}+\bm{\rho}(t)$ where $g(t)=\|{\mathbf{w}}_{t}\|_{2}\rightarrow\infty$, $\bm{\rho}(t)\in\mathbb{R}^{d}$, and $\frac{1}{g(t)}\bm{\rho}(t)\rightarrow\mathbf{0}$. Then, the mini-batch gradient can be represented by
@@
Then we move on to the case of $\ell=\ell_{\text{log}}$. This kind of extension is possible since the logistic loss has a similar tail behavior of the exponential loss, following the line of Soudry et al. [2018]. We adopt the same notation with previous part, and we decompose dominant and residual terms as follows:
@@ Appendix G Missing Proofs in Section 5
Our proof of Theorem 5.1 builds on standard techniques from the analysis of the implicit bias of normalized steepest descent on linearly separable data [Gunasekar et al., 2018a, Zhang et al., 2024a, Fan et al., 2025]. The most closely related result is due to Fan et al. [2025], who showed that full-batch Signum converges in direction to the maximum $\ell_{\infty}$-margin solution. Theorem 5.1 extends this result to the mini-batch setting, establishing that the mini-batch variant of Inc-Signum (Algorithm 4) also converges in direction to the maximum $\ell_{\infty}$-margin solution, provided the momentum parameter is chosen sufficiently close to $1$.
@@
We note that as an extreme case, Inc-Signum with $\beta=0$ and batch size $1$ (i.e., SignSGD) has a simple implicit bias: its iterates converge in direction to $\sum_{i\in[N]}\mathrm{sign}({\mathbf{x}}_{i})$, which corresponds to neither the $\ell_{2}$- nor the $\ell_{\infty}$-max-margin solution.
@@ I.1 Proxy Function
This lemma (or a similar variant) is proved in Zhang et al. [2024a] and Fan et al. [2025]. Below, we provide a proof for completeness.
@@
    For exponential loss, $\frac{{\mathcal{G}}({\mathbf{w}})}{\mathcal{L}({\mathbf{w}})}=1$. For logistic loss, the lower bound $\frac{{\mathcal{G}}({\mathbf{w}})}{\mathcal{L}({\mathbf{w}})}\geq 1-\frac{N\mathcal{L}({\mathbf{w}})}{2}$ follows from Zhang et al. [2024a, Lemma C.7]. The upper bound follows from the elementary inequality $\ell^{\prime}_{\log}(z)=\frac{\exp(-z)}{1+\exp(-z)}\leq\log(1+\exp(-z))=\ell_{\log}(z)$ for all $z\in\mathbb{R}$.
@@ I.2 Properties of Loss Functions
Lemma I.2 (Lemma C.4 in Zhang et al. [2024a]).
@@
Lemma I.3 (Lemma C.5 in Zhang et al. [2024a]).
@@
Lemma I.4 (Lemma C.6 in Zhang et al. [2024a]).
@@ I.3 Auxiliary Results
Lemma I.6 (Lemma C.1 in Zhang et al. [2024a]).

### v3

@@ Appendix E Missing Proofs in Section 3
In this section, we provide the omitted proofs in Section 3. We first introduce the proof of Corollary 3.2 describing how SR datasets eliminate coordinate-adaptivity of Inc-Adam. Then, we review previous literature on the limit direction of weighted GD and prove Theorem 3.3.
@@
Given SR data $\{{\mathbf{x}}_{i}\}_{i\in[N]}$, let $x_{i}=|{\mathbf{x}}_{i}[0]|$. Notice that
@@
We now turn to the proof of Theorem 3.3, building upon the foundational work of Ji et al. (2020), who characterized the convergence direction of GD via its regularization path. Subsequent research has extended this characterization to weighted GD, which optimizes the weighted empirical risk $\mathcal{L}_{\mathbf{q}(t)}({\mathbf{w}})=\sum_{i\in[N]}q_{i}(t)\ell({\mathbf{w}}^{\top}{\mathbf{x}}_{i})$. Xu et al. (2021) proved that weighted GD converges to $\ell_{2}$-max-margin direction on the same linear classification task when the weights are fixed during training. This condition was later relaxed by Zhai et al. (2023), who demonstrated that the same convergence guarantee holds provided the weights converge to a limit, i.e., $\exists\lim_{t\rightarrow\infty}\mathbf{q}(t)=\hat{\mathbf{q}}$.
Our setting, however, introduces distinct technical challenges. First, the weights are bounded but not guaranteed to converge. The most relevant existing result is Theorem 7 in Zhai et al. (2023), which establishes the same limit direction but requires the stronger combined assumptions of lower-bounded weights, loss convergence, and directional convergence of the iterates. A further complication in our analysis is an additional error term, $\bm{\epsilon}_{r}$ in Corollary 3.2, which must be carefully controlled. Our fine-grained analysis overcomes these issues by extending the methodology of Ji et al. (2020), enabling us to manage the error term under the sole, weaker assumption of loss convergence.
@@
By introducing $\bm{a}$}-weighted loss, we can regard weighted GD as vanilla GD with respect to weighted loss. To follow the line of Ji et al. (2020), we show that the regularization path converges in direction to $\ell_{2}$-max-margin solution, regardless of the choice of the weight vector $\bm{a}$} if it is bounded by two positive constants, and such convergence is uniform; we can take sufficiently large $B$ to be close the $\ell_{2}$ solution for any ${\bm{a}}\in[c_{1},c_{2}]^{N}$.
Lemma E.2 (Adaptation of Proposition 10 in Ji et al. (2020)).
@@
We first have to show the uniqueness of $\ell_{2}$-max-margin solution. This proof was introduced by Ji et al. (2020, Proposition 10), but we provide it for completeness. Suppose that there exist two distinct unit vectors ${\mathbf{u}}_{1}$ and ${\mathbf{u}}_{2}$ such that both of them achieve the max-margin $\hat{\gamma}$. Take ${\mathbf{u}}_{3}=\frac{{\mathbf{u}}_{1}+{\mathbf{u}}_{2}}{2}$ as a middle point of ${\mathbf{u}}_{1}$ and ${\mathbf{u}}_{2}$. Then we get
@@
Lemma E.3 (Adaptation of Lemma 9 in Ji et al. (2020)).
@@ Appendix F Missing Proofs in Section 4
which implies $\sum_{t=t_{0}}^{\infty}\eta_{t}\|\nabla\mathcal{L}({\mathbf{w}}_{t})\|_{{\mathbf{P}}^{-1}({\mathbf{w}}_{t})}<\infty$. Since $\sum_{t=t_{0}}^{T}\eta_{t}=\infty$, we get $\liminf_{t\rightarrow\infty}\|\nabla\mathcal{L}({\mathbf{w}}_{t})\|_{{\mathbf{P}}^{-1}({\mathbf{w}}_{t})}=0$. From Lemma F.1, we get $\liminf_{t\rightarrow\infty}\|\nabla\mathcal{L}({\mathbf{w}}_{t})\|_{2}=0$, also implying $\liminf_{t\rightarrow\infty}\mathcal{L}({\mathbf{w}}_{t})=0$. Since $\mathcal{L}({\mathbf{w}}_{t})$ is monotonically decreasing, we get $\mathcal{L}({\mathbf{w}}_{t})\rightarrow 0$. ∎
@@
Before we provide a rigorous proof of Lemma 4.5, we first demonstrate its intuitive explanation motivated by Soudry et al. (2018). For simplicity, assume $\ell=\ell_{\exp}$ and let ${\mathbf{w}}_{t}=g(t)\hat{{\mathbf{w}}}+\bm{\rho}(t)$ where $g(t)=\|{\mathbf{w}}_{t}\|_{2}\rightarrow\infty$, $\bm{\rho}(t)\in\mathbb{R}^{d}$, and $\frac{1}{g(t)}\bm{\rho}(t)\rightarrow\mathbf{0}$. Then, the mini-batch gradient can be represented by
@@
Then we move on to the case of $\ell=\ell_{\text{log}}$. This kind of extension is possible since the logistic loss has a similar tail behavior of the exponential loss, following the line of Soudry et al. (2018). We adopt the same notation with previous part, and we decompose dominant and residual terms as follows:
@@ Appendix G Missing Proofs in Section 5
Our proof of Theorem 5.1 builds on standard techniques from the analysis of the implicit bias of normalized steepest descent on linearly separable data (Gunasekar et al., 2018b; Zhang et al., 2024a; Fan et al., 2025). The most closely related result is due to Fan et al. (2025), who showed that full-batch Signum converges in direction to the maximum $\ell_{\infty}$-margin solution. Theorem 5.1 extends this result to the mini-batch setting, establishing that the mini-batch variant of Inc-Signum (Algorithm 4) also converges in direction to the maximum $\ell_{\infty}$-margin solution, provided the momentum parameter is chosen sufficiently close to $1$.
@@
We note that as an extreme case, Inc-Signum with $\beta=0$ and a batch size of $1$ (i.e., SignSGD) has a simple implicit bias: its iterates converge in direction to $\sum_{i\in[N]}\mathrm{sign}({\mathbf{x}}_{i})$, which corresponds to neither the $\ell_{2}$- nor the $\ell_{\infty}$-max-margin solution.
@@ I.1 Proxy Function
This lemma (or a similar variant) is proved in Zhang et al. (2024a) and Fan et al. (2025). Below, we provide a proof for completeness.
@@
    For exponential loss, $\frac{{\mathcal{G}}({\mathbf{w}})}{\mathcal{L}({\mathbf{w}})}=1$. For logistic loss, the lower bound $\frac{{\mathcal{G}}({\mathbf{w}})}{\mathcal{L}({\mathbf{w}})}\geq 1-\frac{N\mathcal{L}({\mathbf{w}})}{2}$ follows from Zhang et al. (2024a, Lemma C.7). The upper bound follows from the elementary inequality $\ell^{\prime}_{\log}(z)=\frac{\exp(-z)}{1+\exp(-z)}\leq\log(1+\exp(-z))=\ell_{\log}(z)$ for all $z\in\mathbb{R}$.
@@ I.2 Properties of Loss Functions
Lemma I.2 (Lemma C.4 in Zhang et al. (2024a)).
@@
Lemma I.3 (Lemma C.5 in Zhang et al. (2024a)).
@@
Lemma I.4 (Lemma C.6 in Zhang et al. (2024a)).
@@ I.3 Auxiliary Results
Lemma I.6 (Lemma C.1 in Zhang et al. (2024a)).